The financial statements of Snapit Company are given below.

Snapit Company
Income Statement (2009)

Sales $ 4,000,000
Cost of goods sold 3,040,000
Gross profit 960,000
Selling & administrative expenses 430,000
Operating profit 530,000
Interest expense 160,000
Income before tax 370,000
Tax expense 148,000
Net income $ 222,000
Balance Sheet 2009 2008
Cash $ 60,000 $ 50,000
Accounts receivable 550,000 500,000
Inventory 690,000 620,000
Total current assets $ 1,300,000 $ 1,170,000
Fixed assets 1,300,000 1,230,000
Total assets $ 2,600,000 $ 2,400,000
Accounts Payable $ 270,000 $ 250,000
Bank loan 580,000 500,000
Total current liabilities $ 850,000 $ 750,000
Bond payable 900,000 1,000,000
Total liabilities $ 1,750,000 $ 1,750,000
Common stock (130,000 shares) $ 250,000 $ 250,000
Retained earnings 600,000 400,000
Total liabilities & equity $ 2,600,000 $ 2,400,000

Note: The common shares are trading in the stock market for $100 each.

Refer to the financial statements of Snapit Company. The firm's asset turnover ratio for 2009 is

Multiple Choice

  • 3.16.

  • 3.31.

  • 1.60.

  • 4.64.
Answer #1

Answer is 1.6

The firm's asset turnover ratio =Net Sales/Average Total asset

Average Total asset=(2600000+2400000)/2=$2,500,000

firm's asset turnover ratio=4,000,000/2,500,000=1.60